OpenStudy (anonymous):

Find the slope intercept form x intercept-2 y intercept- 3/5

OpenStudy (anonymous):

2x-y 3/5 ???

OpenStudy (anonymous):

would that be the equation its written in?

OpenStudy (anonymous):

slope intercept form is of the form y=mx+b where m is the slope and b is the y intercept

OpenStudy (anonymous):

2x-y + 3 =0 y=m(x)3 y=2x+ 3 2x-y+ 3 = 0??

OpenStudy (anonymous):

is the x intercept = -2 and the y intercept = -3/5 ?

OpenStudy (anonymous):

I'm sorry when i originally asked the question they werent negatives… that was just a dash sign